# The Unbounded Denominators Conjecture

### v2

@@ 5.2. Geometry of $\Gamma_{N}$ and a uniform growth estimate of $F_{N}$
Our second aim in the present § 5 is the uniform supremum growth estimate Lemma 5.2.17 of the universal covering map $F_{N}:D(0,1)\to\mathbf{C}\smallsetminus\mu_{N}$ near the boundary, as both the circle $|x|=r$ and the level $N$ vary. This result is subsumed by the more precise bound of Kraus and Roth [KR16, Theorems 1.2 and 1.10]; our treatment is self-contained. In § 6 we will refine this supremum growth bound (uniformly exponential in $\frac{N}{1-r}$) to an integrated growth bound (uniformly linear in $\frac{N}{1-r}$).
@@
Consider the Dirichlet domain $\Omega_{N}\subset D(0,1)$ associated to $\Gamma_{N}$ around $x=0$. More precisely, we can describe $\Omega_{N}$ as the region:
$\{z\in D(0,1)\mid d(gx,0)\geq d(x,0),\quad g,g^{-1}\in\{r_{N}^{k}tr_{N}^{-k}\},k=0,1,\ldots,N-1\}$
The region in $D(0,1)$ such that $d(gx,0)\geq d(x,0)$ and $d(g^{-1}x,0)\geq d(x,0)$ for $g=r_{N}^{k}tr_{N}^{-k}$ is the region bounded by two geodesics starting at $\zeta_{N}^{k}$ going in opposite directions and intersecting the boundary at $\zeta_{N}^{k}e^{\pm i\theta}$ where $c_{N}=2\cot(\theta/2)$. There are exactly $2N$ such arcs corresponding to the $N$ generators and their inverses. In particular, if $\theta<\pi/N$ is too small, the fundamental region will have infinite volume, whereas if $\theta>\pi/N$ is too big, then the Dirichlet domain will only contain at most $N$ cusps. Since $\mathbf{H}/\widetilde{\Gamma}_{N}$ has $N+1$ cusps and $\widetilde{\Gamma}_{N}$ has finite covolume, we must have $c_{N}=2\cot(\pi/2N)$ and these geodesics intersecting at $\zeta_{N}^{k+1/2}$, $k=0,\ldots,N-1$. ∎
@@
Recall our normalization of $\widetilde{F}_{N}$ that $\widetilde{F}_{N}(i)=0$ and $\widetilde{F}_{N}(i\infty)=1$. By Corollary 5.2.4, $\tau=\pm\cot(\pi/2N)$ (in the same $\widetilde{\Phi}_{N}$-orbit) is the other cusp of $\widetilde{\Phi}_{N}$ and thus $\widetilde{F}_{N}(\pm\cot(\pi/2N))=\infty$. Moreover, both sides of (5.2.6) are uniformizers of $\mathbf{H}/\widetilde{\Phi}_{N}$ which take the value 0 at the cusp $\cot(\pi/2N)$ and $\infty$ at the cusp $i\infty$. This specifies them uniquely up to $z\mapsto\lambda z$ scalings. However, this last ambiguity is removed by noting that both sides are $1$ at $\tau=i$. ∎
@@
The group $\widetilde{\Phi}_{N}$ is contained with index two in the larger group $\widetilde{\Psi}_{N}=\langle\widetilde{s},\widetilde{t}_{N}\rangle$. The group $\widetilde{\Psi}_{N}$ has a fundamental domain consisting of the points $0,1,\zeta_{N}^{1/2}$. But this is none other than a hyperbolic triangle with angles $\{\alpha,\beta,\gamma\}=\{\pi/N,0,0\}$, whose conformal mapping from $\mathbf{H}$ is given by Schwarz triangle functions (see for instance [Car54, §404 on p. 185]). This suggests that $\widetilde{F}_{N}$ should directly be related to Schwarz triangle functions, which leads to a direct description of the inverse functions $\psi_{N}$ and $\varphi_{N}$ in terms of hypergeometric functions in Lemma 5.1.14.
In order to study the behavior of $\widetilde{F}_{N}$ near $z=1$, we use the explicit formula for its inverse $\psi_{N}$ given in Lemma 5.1.14. The following lemma gives the asymptotic of the function $s_{N}$ used in formula of $\psi_{N}$.
@@
Fix a real constant $M_{0}>0$. For $M\geq M_{0}$ and $|z|<e^{-MN}$, we have the uniform estimate:
(5.2.9) $\left|\frac{s_{N}(1-z)}{\gamma_{N}}-(1-z)^{1/N}\frac{-\log z+2\gamma-2\psi(1/2+1/2N)}{-\log z+2\gamma-2\psi(1/2-1/2N)}\right|\ll\frac{|z|}{N}$
where the implied estimates depend on $M_{0}$ but not on $N,M$ and $\psi$ is the digamma function as in Lemma 5.1.18.
@@
By Lemma 5.1.18, for $a=1/2\pm 1/2N$ and $|z|<1$, we have the following equality:
@@
We first prove that the coefficients in this power series are uniformly bounded. Since $|a|<1$, we have $|(a)_{k}|/k!<1$; moreover by the properties of $\psi$-functions (see for instance [AS92, 6.3.5, 6.3.14]), we have $\psi(z)$ is negative and strictly increasing for $0<z<1$ and $|\psi(k)-\psi(k+a)|>|\psi(k+1)-\psi(k+1+a)|$; thus $|\psi(k)-\psi(k+a)|$ is maximized when $k=1$ and a=1/21/4a=1/2-1/4. This immediately leads to the uniform estimates:
$\left|\sum_{k=1}^{\infty}\frac{(a)_{k}(a)_{k}}{k!^{2}}z^{k}\right|,\quad\left|\sum_{k=1}^{\infty}\frac{(a)_{k}(a)_{k}}{k!^{2}}z^{k}\left(2(\psi(k+1)-\psi(k+a))\right)\right|\ll|z|<e^{-MN}$
For $|z|<e^{-MN}$, we also have $\mathrm{Re}(\log z)\leq-MN$, and so in particular $|\log z|\geq MN$ regardless of the branch of logarithm. Combined with Lemma 5.1.14 and (5.2.10), this leads to the estimate
$\frac{s_{N}(1-z)}{\gamma_{N}(1-z)^{1/N}}=\frac{-\log(z)+2\psi(1)-2\psi(1/2+1/2N)+O(z)}{-\log(z)+2\psi(1)-2\psi(1/2-1/2N)+O(z)}$
where the implied constants are uniform in $N$, from which the result follows (using that $|\log z|\gg N$). ∎
@@
$\tau=\widetilde{\psi}_{N}(1-z),\qquad\text{for }|z|<e^{-M(1-\varepsilon_{0})N}.$
@@
(5.2.13) $\tau=i\cdot\frac{1+\psi_{N}(1-z)}{1-\psi_{N}(1-z)}=i\cdot\frac{\gamma_{N}+s_{N}((1-z)^{N})}{\gamma_{N}-s_{N}((1-z)^{N})}.$
Write $(1-z)^{N}=1-Z$, then the same estimate as above implies $|Z|<e^{-M(1-2\varepsilon_{0})N}$ for sufficiently large $N$. Thus we reduce the lemma to the estimate of
(5.2.14) $\tau=i\cdot\frac{\gamma_{N}+s_{N}(1-Z)}{\gamma_{N}-s_{N}(1-Z)},\quad|Z|<e^{-M(1-2\varepsilon_{0})N}.$
@@
$\tau=\frac{i\cot(\pi/2N)}{\pi}\left(2\gamma-\log Z-\psi(1/2-1/2N)-\psi(1/2+1/2N)\right)+O(1),$
where the implicit constant only depends on $M_{0}$ and $\varepsilon_{0}$. The imaginary part of this does not depend on the choice of branch of $\log Z$ and indeed only depends on $|Z|$, and we deduce with this approximation that
$\mathrm{im}(\tau)\geq\frac{\cot(\pi/2N)}{\pi}\left(2\gamma+NM(1-2\varepsilon_{0})-\psi(1/2-1/2N)-\psi(1/2+1/2N)\right)+O(1).$
@@
We close this section by using Lemma 5.2.11 to derive a coarse yet fairly uniform upper bound on $\sup_{|x|=r}\log|F_{N}(x)|$. Although not best-possible, it is enough as an input for the logarithmic error term in the Nevanlinna theory estimate in §6.
@@
Set $S(M,N):=\{x\in D(0,1)\,:\,|F_{N}^{N}(x)-1|<e^{-MN}\}$ and set $M=\frac{N}{1-r}$. Since $M\geq 2$, we may take $M_{0}=2$ and $\epsilon=1/2$ in Lemma 5.2.11 and conclude that for $N\gg 1$ (with absolute constant here), we have
@@
By Shimizu’s Lemma (see, for example, [EGM98, Theorem 3.1]) and Proposition 5.2.1, we have  $2|c|\cot(\pi/2N)\geq 1$ for all $\widetilde{\gamma}=\left(\begin{matrix}a&b\\ c&d\end{matrix}\right)\in\widetilde{\Gamma}_{N}$. Thus by Lemma 5.2.15,
$E(\gamma,D)\leq 2D^{-1}c^{-2}\ll N^{-2}N^{-1}(1-r)N^{2}=\frac{1-r}{N},$
@@
By Lemma 5.2.5, the set $\{x\in D(0,1)\,:\,|F_{N}(x)|>e^{M}+1\}$ is contained in $sS(M,N)$, which is contained in $D(0,1)\smallsetminus\overline{D(0,r)}$ by the above argument for $N\gg 1$. Thus we conclude that
$\sup_{|x|=r}\log|F_{N}(x)|\leq\log(e^{M}+1)\ll M=\frac{N}{1-r}.\qed$
@@ 6.1.2. The lemma on the logarithmic derivative
and proved that it is essentially best-possible in form apart for the value of the free numerical constant $5.8501$ (that has since been somewhat further reduced in the literature, see Benbourenane–Korhonen [BK01]). The paper of Hinkkanen [Hin92] and the book of Cherry and Ye [CY01] discuss the implications to the structure of the error term in Nevanlinna’s second main theorem, mirroring Osgood and Vojta’s dictionary to Diophantine approximation and comparing to Lang’s conjecture modeled on Khinchin’s theorem.
@@ 6.2.2. Reduction to a logarithmic derivative
The idea then is that the piece $f^{\prime}/f$ in the decomposition 6.2.5 is small on average over circles by the lemma on the logarithmic derivative (Corollary 6.2.6 below), while—again by the lemma on the logarithmic derivative, in Corollary 6.2.8 below—the characteristic functions of $p\circ F_{N}$ and $F_{N}/F_{N}^{\prime}$ are equal respectively to $NT(r,F_{N})$ and $T(r,F_{N})$ up to a small error.
@@ 6.2.3. Two corollaries of the lemma on the logarithmic derivative
(6.2.7) $m\Big(r,\frac{f^{\prime}}{f}\Big)\ll\sup_{|z|=(1+r)/2}\log^{+}{\log{|F_{N}|}}+\log^{+}{\frac{N}{1-r}},$
@@
$m\Big(r,\frac{f^{\prime}}{f}\Big)\ll\log^{+}{T\Big(\frac{1+r}{2},f\Big)}+\log^{+}{\frac{e}{1-r}}$
$=\log^{+}{m\Big(\frac{1+r}{2},1-F_{N}^{N}\Big)}+\log^{+}{\frac{e}{1-r}}$
$\ll\log^{+}{m\Big(\frac{1+r}{2},F_{N}^{N}\Big)}+\log^{+}{\frac{e}{1-r}}$
$\ll\log^{+}{m\Big(\frac{1+r}{2},F_{N}\Big)}+\log^{+}{\frac{N}{1-r}}$
$\ll\sup_{|z|=(1+r)/2}\log^{+}{\log{|F_{N}|}}+\log^{+}{\frac{N}{1-r}},$

### v3

@@ 5.2. Geometry of $\Gamma_{N}$ and a uniform growth estimate of $F_{N}$
Our second aim in the present § 5 is the uniform supremum growth estimate Lemma 5.2.17 of the universal covering map $F_{N}:D(0,1)\to\mathbf{C}\smallsetminus\mu_{N}$ near the boundary, as both the circle $|z|=r$ and the level $N$ vary. This result is subsumed by the more precise bound of Kraus and Roth [KR16, Theorems 1.2 and 1.10]; our treatment is self-contained. In § 6 we will refine this supremum growth bound (uniformly exponential in $\frac{N}{1-r}$) to an integrated growth bound (uniformly linear in $\frac{N}{1-r}$).
@@
Consider the Dirichlet domain $\Omega_{N}\subset D(0,1)$ associated to $\Gamma_{N}$ around $z=0$. More precisely, we can describe $\Omega_{N}$ as the region
$\{z\in D(0,1)\,:\,d(gz,0)\geq d(z,0)\quad\forall g,g^{-1}\in\{r_{N}^{k}\,t\,r_{N}^{-k}\},k=0,1,\ldots,N-1\}.$
Here, $d$ is the hyperbolic distance in $D(0,1)$. The region in $D(0,1)$ such that $d(gz,0)\geq d(z,0)$ and $d(g^{-1}z,0)\geq d(z,0)$ for $g=r_{N}^{k}\cdot t\cdot r_{N}^{-k}$ is the region bounded by two geodesics starting at $\zeta_{N}^{k}$ going in opposite directions and intersecting the boundary at $\zeta_{N}^{k}e^{\pm i\theta}$ where $c_{N}=2\cot(\theta/2)$. There are exactly $2N$ such arcs corresponding to the $N$ generators and their inverses. In particular, if $\theta<\pi/N$ is too small, the fundamental region will have infinite volume, whereas if $\theta>\pi/N$ is too big, then the Dirichlet domain will only contain at most $N$ cusps. Since $\mathbf{H}/\widetilde{\Gamma}_{N}$ has $N+1$ cusps and $\widetilde{\Gamma}_{N}$ has finite covolume, we must have $c_{N}=2\cot(\pi/2N)$ and these geodesics intersecting at $\zeta_{N}^{k+1/2}$, $k=0,\ldots,N-1$. ∎
@@
Recall our normalization of $\widetilde{F}_{N}$ that $\widetilde{F}_{N}(i)=0$ and $\widetilde{F}_{N}(i\infty)=1$. By Corollary 5.2.4, $\tau=\pm\cot(\pi/2N)$ (in the same $\widetilde{\Phi}_{N}$-orbit) is the other cusp of $\widetilde{\Phi}_{N}$ and thus $\widetilde{F}_{N}(\pm\cot(\pi/2N))=\infty$. Moreover, both sides of (5.2.6) are uniformizers of $\mathbf{H}/\widetilde{\Phi}_{N}$ which take the value 0 at the cusp $\cot(\pi/2N)$ and $\infty$ at the cusp $i\infty$. This specifies them uniquely up to $x\mapsto\lambda x$ scalings. However, this last ambiguity is removed by noting that both sides are $1$ at $\tau=i$. ∎
@@
The group $\widetilde{\Phi}_{N}$ is contained with index two in the larger group $\widetilde{\Psi}_{N}=\langle\widetilde{s},\widetilde{t}_{N}\rangle$. The group $\widetilde{\Psi}_{N}$ has a fundamental domain consisting of the points $0,1,\zeta_{N}^{1/2}$. But this is none other than a hyperbolic triangle with angles $\{\alpha,\beta,\gamma\}=\{\pi/N,0,0\}$, whose conformal mapping from $\mathbf{H}$ is given by Schwarz triangle functions (see, for instance, [Car54, §404 on page 185]). This suggests that $\widetilde{F}_{N}$ should directly be related to Schwarz triangle functions, which leads to a direct description of the inverse functions $\psi_{N}$ and $\varphi_{N}$ in terms of hypergeometric functions in Lemma 5.1.14.
In order to study the behavior of $\widetilde{F}_{N}$ near $x=1$, we use the explicit formula for its inverse $\psi_{N}$ given in Lemma 5.1.14. The following lemma gives the asymptotic of the function $s_{N}$ used in formula of $\psi_{N}$.
@@
Fix a real constant $M_{0}>0$. For $M\geq M_{0}$ and $|x|<e^{-MN}$, we have the uniform estimate:
(5.2.9) $\left|\frac{s_{N}(1-x)}{\gamma_{N}}-(1-x)^{1/N}\frac{-\log x+2\gamma-2\psi(1/2+1/2N)}{-\log x+2\gamma-2\psi(1/2-1/2N)}\right|\ll\frac{|x|}{N},$
where $\psi$ is the digamma function as in Lemma 5.1.18, and the implicit coefficient depends on $M_{0}$ but not on $N,M$.
@@
By Lemma 5.1.18, we have for $a=1/2\pm 1/2N$ and $|x|<1$ the following equality:
@@
We first prove that the coefficients in this power series are uniformly bounded. Since $|a|<1$, we have $|(a)_{k}|/k!<1$. Basic properties of the digamma function (cf. [AS92, 6.3.5, 6.3.14]) show that $\psi(x)$ is negative and strictly increasing for $0<x<1$, and $|\psi(k)-\psi(k+a)|>|\psi(k+1)-\psi(k+1+a)|$. Hence, $|\psi(k)-\psi(k+a)|$ is maximized when $k=1$ and a=1/21/4a=1/2-1/4. This immediately leads to the uniform estimates
$\left|\sum_{k=1}^{\infty}\frac{(a)_{k}(a)_{k}}{k!^{2}}x^{k}\right|,\quad\left|\sum_{k=1}^{\infty}\frac{(a)_{k}(a)_{k}}{k!^{2}}x^{k}\left(2(\psi(k+1)-\psi(k+a))\right)\right|\ll|x|<e^{-MN}$
For $|x|<e^{-MN}$, we also have $\Re(\log{x})\leq-MN$, and so in particular $|\log{x}|\geq MN$ regardless of the branch of logarithm. Combined with Lemma 5.1.14 and (5.2.10), this leads to the estimate
$\frac{s_{N}(1-x)}{\gamma_{N}(1-x)^{1/N}}=\frac{-\log{x}+2\psi(1)-2\psi(1/2+1/2N)+O(x)}{-\log{x}+2\psi(1)-2\psi(1/2-1/2N)+O(x)}$
where the implied constants are uniform in $N$, from which the result follows (using that $|\log{x}|\gg N$). ∎
@@
$\tau=\widetilde{\psi}_{N}(1-x),\qquad\text{for }|x|<e^{-M(1-\varepsilon_{0})N}.$
@@
(5.2.13) $\tau=i\cdot\frac{1+\psi_{N}(1-x)}{1-\psi_{N}(1-x)}=i\cdot\frac{\gamma_{N}+s_{N}((1-x)^{N})}{\gamma_{N}-s_{N}((1-x)^{N})}.$
Writing $1-X:=(1-x)^{N}$, then the same estimate as above implies $|X|<e^{-M(1-2\varepsilon_{0})N}$ for sufficiently large $N$. Thus we reduce the lemma to the estimate of
(5.2.14) $\tau=i\cdot\frac{\gamma_{N}+s_{N}(1-X)}{\gamma_{N}-s_{N}(1-X)},\quad|X|<e^{-M(1-2\varepsilon_{0})N}.$
@@
$\tau=\frac{i\cot(\pi/2N)}{\pi}\left(2\gamma-\log{X}-\psi(1/2-1/2N)-\psi(1/2+1/2N)\right)+O(1),$
where the implicit constant only depends on $M_{0}$ and $\varepsilon_{0}$. The imaginary part of this does not depend on the choice of branch of $\log{X}$ and indeed only depends on $|X|$, and we deduce with this approximation that
$\Im(\tau)\geq\frac{\cot(\pi/2N)}{\pi}\left(2\gamma+NM(1-2\varepsilon_{0})-\psi(1/2-1/2N)-\psi(1/2+1/2N)\right)+O(1).$
@@
We close this section by using Lemma 5.2.11 to derive a coarse yet fairly uniform upper bound on $\sup_{|z|=r}\log|F_{N}(z)|$. Although not best-possible, it is enough as an input for the logarithmic error term in the Nevanlinna theory estimate in §6.
@@
Set $S(M,N):=\{z\in D(0,1)\,:\,|F_{N}^{N}(z)-1|<e^{-MN}\}$ and $M:=\frac{N}{1-r}$. Since $M\geq 2$, we may take $M_{0}=2$ and $\epsilon=1/2$ in Lemma 5.2.11 and conclude that for $N\gg 1$ (with absolute constant here), we have
@@
By Shimizu’s Lemma (see, for example, [EGM98, Theorem 3.1]) and Proposition 5.2.1, we have  $2(|a|+|c|)\cot(\pi/2N)\geq 1$ for all $\widetilde{\gamma}=\left(\begin{matrix}a&b\\ c&d\end{matrix}\right)\in\widetilde{\Gamma}_{N}$. Thus by Lemma 5.2.15,
$E(\gamma,D)\leq 2D^{-1}(a^{2}+c^{2})\ll N^{-2}N^{-1}(1-r)N^{2}=\frac{1-r}{N},$
@@
By Lemma 5.2.5, the set $\{z\in D(0,1)\,:\,|F_{N}(z)|>e^{M}+1\}$ is contained in $sS(M,N)$, which is contained in $D(0,1)\smallsetminus\overline{D(0,r)}$ by the above argument for $N\gg 1$. Thus we conclude that
$\sup_{|z|=r}\log|F_{N}(z)|\leq\log(e^{M}+1)\ll M=\frac{N}{1-r}.\qed$
@@ 6.1.2. The lemma on the logarithmic derivative
and proved that it is essentially best-possible in form apart for the value of the free numerical constant $5.8501$ (that has since been somewhat further reduced in the literature, see Benbourenane–Korhonen [BK01]). The paper of Hinkkanen [Hin92] and the books of Cherry–Ye [CY01] and Ru [Ru21] discuss the implications to the structure of the error term in Nevanlinna’s second main theorem, mirroring Osgood and Vojta’s dictionary to Diophantine approximation and comparing to Lang’s conjecture modeled on Khinchin’s theorem.
@@ 6.2.2. Reduction to a logarithmic derivative
The idea then is that the piece $f^{\prime}/f$ in the decomposition (6.2.5) is small on average over circles by the lemma on the logarithmic derivative (Corollary 6.2.6 below), while—again by the lemma on the logarithmic derivative, in Corollary 6.2.8 below—the characteristic functions of $p\circ F_{N}$ and $F_{N}/F_{N}^{\prime}$ are equal respectively to $NT(r,F_{N})$ and $T(r,F_{N})$ up to a small error.
@@ 6.2.3. Two corollaries of the lemma on the logarithmic derivative
(6.2.7) $m\Big(r,\frac{f^{\prime}}{f}\Big)\ll\sup_{|z|=(1+r)/2}\log^{+}{\log{|F_{N}|}}+\log{\frac{N}{1-r}}.$
@@
$m\Big(r,\frac{f^{\prime}}{f}\Big)\ll\log^{+}{T\Big(\frac{1+r}{2},f\Big)}+\log{\frac{e}{1-r}}$
$=\log^{+}{m\Big(\frac{1+r}{2},1-F_{N}^{N}\Big)}+\log{\frac{e}{1-r}}$
$\ll\log^{+}{m\Big(\frac{1+r}{2},F_{N}^{N}\Big)}+\log{\frac{e}{1-r}}$
$\ll\log^{+}{m\Big(\frac{1+r}{2},F_{N}\Big)}+\log{\frac{N}{1-r}}$
$\ll\sup_{|z|=(1+r)/2}\log^{+}{\log{|F_{N}|}}+\log{\frac{N}{1-r}},$